OpenStudy (anonymous):

product of two nos is 21 and difference is 4. then ratio of sum of their cubes to diff. of their cubes is?

OpenStudy (bahrom7893):

xy = 21 x-y = 4 (x^3+y^3)/(x^3-y^3) = ?

OpenStudy (bahrom7893):

solve that i think all u need was setting up help..

OpenStudy (anonymous):

got it thanks

OpenStudy (bahrom7893):

np..

OpenStudy (anonymous):

here is a way to do this problem without actually solving for x and y, just incase you want to see another solution than the one you already have :)

OpenStudy (bahrom7893):

dang joe.. u always improvise! I love that!